# Human causal perception in a cube-stacking task

### v1

@@ 2.1 Task
In our study the participants are tasked in a binary forced choice task to assess whether a stack of two unit-cubes collapses. The specific instruction was to judge, whether the top cube touches the floor if one were to build the tower in real life. An example of what is presented to the participant is shown in fig. 1. It depicts three images showing the respective stack of cubes from three different perspectives. To give an assessment, the participant has to select one of the options placed directly below the images and submit via pressing the next-button. After submitting a trial the next configuration is shown, so there is no time restriction for annotation. The stimuli shown are the same for all participants, however, in a randomized order.
@@ 2.2 Data
In fig. 2 you an see the positions of the geometrical center [^2] Which coincides with the center of mass in this experiment of the upper cube projected into the coordinate system of the lower cube. Note that the positions are first projected into the lower left corner of the cube and then mirrored along the symmetry-axes of a cube in the 2D plane. The gray square depicts the outlines of the lower cube. As made visible the stack remains stable, as long as the position is within bounds of the lower cube.
@@ 3 Results and discussion
For our analysis we approximated the decision surface based on the displacement of the cubes using a multi layer perceptron with 2 hidden layers of 10 neurons each. Figure 3 shows examples for the approximation. The dotted, black line denotes the outlines of the lower cube.
Our results indicate that the perceived area of support forms a rotated square. Figure 3(a) and fig. 3(b) show examples for this. As a result, most participants classify stable configurations towards the corners of the lower cube as unstable. Interestingly, the majority of participants who correctly assess the corner configurations as stable, make errors towards the center of the respective border, again forming a shape similar to a rotated square or circle. An example for this is shown in fig. 3(c). Besides that, we observed that in almost every participant the probability of assessing a stable configuration as unstable grew anti-proportionally to the distance to the nearest corner. This also includes the bias towards safety reported by other works (e.g. [4, 3, 5]). See appendix A for visualizations of all participants. This includes the ground truth, an idealized expected participant using a safety margin as heuristic and an idealized participant with a rotated square heuristic. Naturally the participants differed in safety measures which translates into different sizes of the rectangles/shapes.

### v2

@@ 2.1 Task
In our study the participants are tasked in a binary forced choice task to assess whether a stack of two unit-cubes collapses. The specific instruction was to judge, whether the top cube touches the floor if one were to build the tower in real life. An example of what is presented to the participant is shown in fig.˜1. It depicts three images showing the respective stack of cubes from three different perspectives. To give an assessment, the participant has to select one of the options placed directly below the images and submit via pressing the next-button. After submitting a trial the next configuration is shown, so there is no time restriction for annotation. The stimuli shown are the same for all participants, however, in a randomized order.
@@ 2.2 Data
In fig.˜2 you an see the positions of the geometrical center [^2] Which coincides with the center of mass in this experiment of the upper cube projected into the coordinate system of the lower cube. Note that the positions are first projected into the lower left corner of the cube and then mirrored along the symmetry-axes of a cube in the 2D plane. The gray square depicts the outlines of the lower cube. As made visible the stack remains stable, as long as the position is within bounds of the lower cube.
@@ 3 Results and discussion
For our analysis we approximated the decision surface based on the displacement of the cubes using a multi layer perceptron with 2 hidden layers of 10 neurons each. Figure˜3 shows examples for the approximation. The dotted, black line denotes the outlines of the lower cube.
Our results indicate that the perceived area of support forms a rotated square. Figure˜3(a) and fig.˜3(b) show examples for this. As a result, most participants classify stable configurations towards the corners of the lower cube as unstable. Interestingly, the majority of participants who correctly assess the corner configurations as stable, make errors towards the center of the respective border, again forming a shape similar to a rotated square or circle. An example for this is shown in fig.˜3(c). Besides that, we observed that in almost every participant the probability of assessing a stable configuration as unstable grew anti-proportionally to the distance to the nearest corner. This also includes the bias towards safety reported by other works (e.g. [4, 3, 5]). See appendix˜A for visualizations of all participants. This includes the ground truth, an idealized expected participant using a safety margin as heuristic and an idealized participant with a rotated square heuristic. Naturally the participants differed in safety measures which translates into different sizes of the rectangles/shapes.